# RadZero: Similarity-Based Cross-Attention for Explainable Vision-Language Alignment in Radiology with Zero-Shot Multi-Task Capability

### v1

@@ 4.1 Training dataset
We train our model using the MIMIC-CXR dataset for vision-language alignment. MIMIC-CXR comprises 377,110 CXR images from 227,835 radiographic studies involving 65,379 patients. Each study includes a radiology report and one or more CXR images in either frontal or lateral views. Images are sourced from MIMIC-CXR-JPG [17], and only the findings and impression sections of reports are extracted using the official codebase [^1] https://github.com/MIT-LCP/mimic-cxr. All view positions are considered, and the official dataset split is followed. As described in Section 3.1, finding-sentence extraction is applied, with each study containing an average of 6.45 such sentences. Studies without extracted finding-sentences are discarded, resulting in 352,875 training images and 2,852 for validation.
@@ 4.2 Test datasets
Open-I [7] dataset consists of 3,851 radiology reports and 7,470 CXR images, with multi-label classification annotations for 18 diseases. PadChest [4] consists of 160,868 CXR images collected from 67,000 patients and exhibits a long-tailed distribution with a total of 192 labels. Following [19], we use 39,053 samples annotated by board-certified radiologists. Additionally, PadChest20, introduced in [19], serves as a test set for rare disease evaluation, consisting of 20 classes with fewer than 10 samples each. ChestXray14 [33] provide official test set with 22,433 images and corresponding labels for 14 diseases. CheXpert [13] provide official official test set consists of 500 patients’ images annotated by five board-certified radiologists. Following [19], we perform classification evaluation on five observations: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. ChestXDet10 [21] is a subset of ChestXray14, consisting of 542 images with bounding box annotations for 10 diseases in the official test set. SIIM [35] pneumothorax dataset contains manually annotated segmentation masks for CXRs, of which 11,582 images are publicly available. We adopt the test split of [32], comprising 1,704 images with 458 positive samples. RSNA [30] pneumonia dataset consists of 29,700 frontal-view radiographs with bounding box annotations indicating evidence of pneumonia. We use the test set of 5,337 images released by [34], of which 1,218 are positive. MS-CXR [2] consists of 1,153 image-phrase-bounding box triplets, with images sourced from MIMIC-CXR. The bounding boxes annotated to specific phrases in the report enable more detailed grounding, referred to as phrase grounding. For a fair comparison on the test set of 167 images released by [6], where each phrase corresponds to only one bounding box, we exclude these test images from the training set described in Sec. 4.1.
@@ 4.4 Implementation details
We adopt XrayDINOv2 [5] as the pre-trained vision encoder, which is trained in a unimodal setting using CXR images based on DINOv2 [25]. While the vision encoder was trained with an image resolution of 224, we increase it to 518 for our experiments. The patch size of $14\times 14$ leads to $37\times 37$ patches, yielding a vision patch length $L$ of 1369. The text encoder is MPNet (“all-mpnet-base-v2”) [28], initialized with pre-trained parameters. The Transformer layers are randomly initialized, with a hidden dimension of 768, matching that of both the vision and text encoders. While the vision encoder remains frozen, all other parameters are trainable. Following [27], the learnable temperature parameter is set to $\log(1/0.07)$. The details of model training can be found in Appendix B. The LLM used for extracting finding-sentences is “Llama-3.3-70B-Instruct” [10].
@@ 5 Results
In Sec. 5.1, we compare RadZero with other SOTA approaches in zero-shot classification, grounding, and segmentation tasks. In Sec. 5.2, we analyze the model’s explainability through cross-modal similarity map analysis. In Sec. 5.3, we validate the potential of RadZero for open-vocabulary semantic segmentation. Due to space limitations, the full ablation study is provided in Appendix C.
  Method Open-I PadChest PadChest20 ChestXray14 CheXpert ChestXDet10 SIIM RSNA
@@
Table 1: Comparison of zero-shot classification performance between RadZero and baseline models on various public CXR datasets. The metric is AUC score, with the best results highlighted in bold.
@@
Table 2: Comparison of zero-shot grounding performance between RadZero and baseline models on ChestXDet10. The metric is pointing game accuracy, with the best results highlighted in bold. Lesion abbreviations can be found in the Appendix A.
@@
Table 3: Comparison of zero-shot phrase grounding performance between RadZero and baseline models on MS-CXR. The metric is pointing game accuracy.
@@
Table 4: Comparison of zero-shot segmentation performance between RadZero and baseline models on RSNA and SIIM. Values marked with * are from Wu et al. [34], while we conduct evaluations for other scores. The best results in zero-shot setting are highlighted in bold.
5.1 Zero-shot evaluation
Classification.
Table 1 compares RadZero with other SOTA models in public test datasets. For the five datasets reported by CARZero [19], we extracted the reported results, while for SIIM and RSNA we independently evaluated two open-source models. Notably, RadZero achieved new SOTA performance on Open-I and PadChest. In PadChest, which is a long-tailed dataset containing 192 classes, we outperformed CARZero by 3.1 points, demonstrating a strong generalization in zero-shot classification. We also observe notable gains in PadChest20 (a subset focusing on rare diseases), suggesting that similarity-based vision-language alignment is especially effective in handling less frequent conditions. For the datasets in which our model missed the first place, MedKLIP performed the best on RSNA, while CARZero excelled on ChestXray14, CheXpert, and ChestXDet10. However, MedKLIP underperformed on the latter datasets, and CARZero underperformed on RSNA. In contrast, RadZero showed results comparable to the top-performing models across all four datasets.
The representative classification metric shown in Figure 1 is the average AUC across all datasets. RadZero established a new SOTA, outperforming CARZero by 1.5 percentage points. This improvement stems from our proposed training strategy, which leverages finding-sentences as minimal semantic units in contrastive learning and incorporates multi-positive training to enhance the diversity of both positive and negative samples per image.
Grounding.
Table 2 presents the results of zero-shot grounding on ChestXDet10. We adopted the pointing game scores reported by CARZero for each disease, except for BioViL-T, which we evaluated using the released model. Based on the average scores across all diseases, RadZero achieved the highest score, surpassing CARZero by a margin of 0.079. Further analysis of the results for each lesion revealed that our model demonstrated the best performance across all classes, with the exception of consolidation. This indicates that our similarity map effectively captures the local similarity between text and visual patch embeddings, irrespective of the disease class. Additionally, our architecture can efficiently train with higher input resolutions, enabling more precise grounding.
Table 3 presents the results of zero-shot phrase grounding. The model’s ability to localize the image region corresponding to a text phrase was measured using pointing game accuracy. We evaluated all baselines using publicly available models. RadZero achieved the highest score of 0.844, showing that it accurately interpreted text phrases. This advancement can be attributed to the training process, which effectively learned the fine-grained similarity between image and text, and the utilization of location information during the finding-sentence extraction process.
Segmentation.
@@ 5.1 Zero-shot evaluation
Table 4 summarizes the segmentation results on SIIM and RSNA. To compare zero-shot performance with supervised models, we fine-tuned the pre-trained MGCA model, with percentage values in parentheses indicating the proportion of training data used.
@@
Among zero-shot models, RadZero achieved the best Dice scores on both SIIM and RSNA. Notably, on SIIM, it showed a 71% improvement of Dice score over CARZero, demonstrating superior segmentation capability. However, the gap in RSNA is smaller, which can be attributed to mismatch in granularity—RSNA’s labels are bounding boxes rather than pixel-wise annotations, which hinders the merit of RadZero’s fine-grained similarity prediction.
RadZero remains competitive even when compared to fine-tuned models. It outperformed MGCA (1%) on both SIIM and RSNA, demonstrating the potential of zero-shot segmentation. While its performance is lower than MGCA (10%) and MGCA (100%), RadZero does not require mask labels during training, allowing it to generalize beyond a fixed vocabulary. This advantage is further demonstrated in our open-vocabulary segmentation results in Sec. 5.3.
@@
SIIM contains fine-grained mask labels, making it suitable for evaluating Pix-AUC scores. RadZero achieved the highest Pix-AUC score, surpassing even MGCA (10%), demonstrating its capability to generate well-calibrated similarity maps that effectively distinguish between positive and negative regions. In contrast, MedKLIP and CARZero, which rely on attention maps, resulted in lower Pix-AUC scores of 0.648 and 0.856, respectively. For a fair comparison, we also evaluated CARZero’s pre-softmax logit maps (CARZero (logits)), which improved results but still fell short of RadZero. CARZero (logits) underperformed even its own original Dice score (0.081 vs. 0.100) despite the optimal threshold search, which aligns with our discussion in Sec. 2.3: dot product-based similarity lacks consistent scaling. In contrast, RadZero’s mask prediction map directly represents text–image pixel-level similarity, allowing low similarity for negative samples. This is reflected in its high Pix-AUC score, further supported by our similarity map analysis in Sec. 5.2.
@@
Figure 5: Visualization of cross-modal similarity maps The CXR images of (a), (b), and (c) are from ChestXDet10, representing normal, fibrosis, and effusion in the right lung, respectively. The similarity probability $\hat{l}$ (top-right corner) between each CXR image and the text prompt (bottom-right corner) can be explained through the visualized similarity map $\hat{M}$.
5.2 Cross-modal similarity map analysis
@@
Figure 5 demonstrates that RadZero effectively aligned visual and textual representations through similarity-based cross-attention. Its outputs, the similarity map $\hat{M}$ and probability $\hat{l}$, offer interpretable visualizations as well as quantitative metrics. For the normal image (Figure 5 (a)), the model assigned low similarity (0.072) to the prompt “There is atelectasis,” with a dark similarity map, indicating no strong alignment. In contrast, the prompts “There is no atelectasis” (0.921) and “The lungs are clear” (0.846) produced high similarities, with bright activations in the lung fields, confirming alignment with normality.
@@
Figure 5 (b) shows fibrosis, and “There is fibrosis” resulted in high similarity (0.970) with strong activations in the affected lung. Conversely, RadZero produced much lower scores (0.093 and 0.014) and darker similarity maps for the two prompts indicating normality, effectively distinguishing between normal and abnormal descriptions.
@@
The key strength of our approach is its ability to differentiate anatomical descriptions. For the right-sided pleural effusion in Figure 5 (c), the model assigned high similarity (0.907) to “There is effusion,” with bright activations in the correct region. Notably, “There is right effusion” (0.946) scored even higher, indicating accurate localization, while “There is left effusion” (0.125) resulted in a much lower score and a dark similarity map, showing that the model correctly distinguishes between the left and right lungs.
@@
Overall, these results highlight the explainability of RadZero. The similarity probability can be validated at the pixel-level, ensuring spatially grounded explanations. Thus, RadZero provides enhanced interpretability by offering a transparent rationale for how conclusions are derived.
@@ 5.3 Open-vocabulary semantic segmentation
Figures 6 and 7 show the results of open-vocabulary semantic segmentation for findings and anatomical regions, respectively. Segmentation mask was obtained for each text prompt by thresholding the similarity map $\hat{M}$. When multiple prompts were identified as positive for a pixel, the prompt with the highest similarity was assigned to it.
Figure 6 shows that the model effectively identified lesion locations based on text prompts. Some regions extended beyond ground truth bounding boxes, clearly indicating room for improvement. Yet, among these incorrect predictions were a few instances where the model captured clinically relevant features that were not explicitly annotated. In Figure 6 (b), RadZero captured a chest tube as similar to “pneumothorax”, which is a reasonable association.
Figure 7 illustrates RadZero’s capability to segment anatomical regions without explicit supervision. Although the model struggles to delineate precise boundaries, it identifies approximate regions, suggesting the ability to infer spatial relationships from text. However, occasional misclassifications, such as identifying ribs as the spine, underscore its limitations and the need for further refinement.
These results demonstrate the potential of RadZero for zero-shot open-vocabulary semantic segmentation, aligning text descriptions with medical images. This approach offers a promising direction for interpretable medical image segmentation, with further advancements anticipated. Additional qualitative results are provided in Appendix E.

### v2

@@ 4.1 Training dataset
We train our model using the MIMIC-CXR dataset for VL alignment. MIMIC-CXR comprises 377,110 CXR images from 227,835 radiographic studies involving 65,379 patients. Each study includes a radiology report and one or more CXR images in either frontal or lateral views. Images are sourced from MIMIC-CXR-JPG [17], and only the findings and impression sections of reports are extracted using the official codebase [^1] https://github.com/MIT-LCP/mimic-cxr. All view positions are considered, and the official dataset split is followed. As described in Section 3.1, finding-sentence extraction is applied, with each study containing an average of 6.45 such sentences. Studies without extracted finding-sentences are discarded, resulting in 352,875 training images and 2,852 for validation.
@@ 4.2 Test datasets
Open-I (OI) [7] contains 3,851 radiology reports and 7,470 CXR images with multi-label annotations for 18 diseases. PadChest (PC) [4] comprises 160,868 CXR images from 67,000 patients, with 192 labels showing a long-tailed distribution. Following [19], we use 39,053 samples annotated by board-certified radiologists. Additionally, PadChest20 (PC20), introduced in [19], serves as a test set for rare disease evaluation, consisting of 20 classes with fewer than 10 samples each. ChestXray14 (CXR14) [34] provides official test set with 22,433 images labeled for 14 diseases. CheXpert (CXP) [13] includes a test set of 500 patients’ images annotated by five board-certified radiologists. Following [19], we evaluate classification on five observations: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. ChestXDet10 (CXD10) [22], a subset of CXR14, contains 542 images with bounding box annotations for 10 diseases in the official test set. SIIM [37] pneumothorax dataset provides segmentation masks for 11,582 CXRs; we adopt the test split from [33], which includes 1,704 images with 458 positives. RSNA [31] pneumonia dataset consists of 29,700 frontal CXRs with bounding box annotations; we use the test set from [36] containing 5,337 images, including 1,218 positives. MS-CXR [2] consists of 1,153 image-phrase-bounding box triplets, with images sourced from MIMIC-CXR. The bounding boxes annotated to specific phrases in the report enable more detailed grounding, referred to as phrase grounding. For fair evaluation on the 167-image test set released by [6], where each phrase maps to a single bounding box, we exclude these images from the training set described in Sec. 4.1.
@@ 4.4 Implementation details
We adopt XrayDINOv2 [5] as the pre-trained vision encoder, which is trained in a unimodal setting using CXR images based on DINOv2 [26]. While the vision encoder was trained with an image resolution of 224, we increase it to 518 for our experiments. The patch size of $14\times 14$ leads to $37\times 37$ patches, yielding a vision patch length $L$ of 1369. The text encoder is MPNet ("all-mpnet-base-v2") [29], initialized with pre-trained parameters. The trainable Transformer layers consist of two randomly initialized layers ($K=2$), with a hidden dimension of 768, matching the embedding size of both the vision and text encoders. While the vision encoder remains frozen, all other parameters are trainable. Following [28], the learnable temperature parameter $\tau$ is initialized to $\log(1/0.07)$. The details of model training can be found in Appendix B. The LLM used for extracting finding-sentences is "Llama-3.3-70B-Instruct" [10], deployed in a private computing environment.
@@ 5 Results
5.1 Zero-shot evaluation
  Method Open-I (OI) PadChest (PC) PadChest20 (PC20) ChestXray14 (CXR14) CheXpert (CXP) ChestXDet10 (CXD10) SIIM RSNA
@@ 5.1 Zero-shot evaluation
Table 1: Zero-shot classification AUROC scores on public CXR datasets. For fair comparison, we also report the results of low-resolution (224×224) version of RadZero.
Classification.
Table 1 compares RadZero with SOTA models on public test datasets. For the five datasets evaluated in CARZero [19], we report their published results. For SIIM and RSNA, we independently evaluated two open-source models. RadZero achieved new SOTA performance on OI and PC, irrespective of image resolution. In the long-tailed PC dataset with 192 classes, RadZero outperformed CARZero by 3.1 percentage points, demonstrating strong generalization in zero-shot classification. Notable gains are also observed in PC20, which focuses on rare diseases, suggesting that VL-CABS is particularly effective for infrequent conditions. On datasets where RadZero failed to rank first, MedKLIP performed best on RSNA, while CARZero led on CXR14, CXP, and CXD10. However, MedKLIP underperformed on the latter datasets, and CARZero underperformed on RSNA. In contrast, RadZero showed results comparable to the top-performing models across all datasets. Interestingly, the lower-resolution RadZero (224px) even outperformed RadZero: potentially due to the pre-trained vision encoder. as discussed in Sec. 5.4.
The representative classification metric shown in Figure 2 is the average AUC across all datasets. RadZero established a new SOTA, outperforming CARZero by 1.5 percentage points, a gain attributable to our training strategy that incorporates multi-positive contrastive learning to enhance the diversity of both positive and negative samples per image.
Grounding.
@@
Table 2: Zero-shot grounding results (pointing game accuracy) on CXD10. Lesion abbreviations can be found in the Appendix E.
@@
Table 3: Zero-shot phrase grounding results (pointing game accuracy) on MS-CXR.
Table 2 presents zero-shot grounding results on CXD10. We adopted the pointing game scores reported by CARZero for all models except BioViL-T, which we evaluated using its released weights. RadZero achieved the highest average score across all diseases, outperforming CARZero by 0.079. Per-lesion analysis showed that RadZero achieved the best performance in all classes except consolidation, indicating that the proposed VL-CABS effectively captures local alignment between text and image patches regardless of disease type. Furthermore, RadZero efficiently supports higher input resolutions, enabling more precise localization.
Table 3 reports zero-shot phrase grounding reports, with pointing game accuracy as the evaluation metric. We evaluated all baselines using publicly available models. RadZero achieved the highest score of 0.844, demonstrating accurate interpretation of text phrases. The strong performance of RadZero (224px) suggests that the gains are largely attributable to the effectiveness of VL-CABS rather than the higher input resolution.
Segmentation.
@@
Table 4: Zero-shot segmentation results. Values with are from [36] and from [21]
@@
Table 4 summarizes zero-shot segmentation results on SIIM and RSNA. To benchmark against supervised models, we fine-tuned MGCA with varying proportions of training data; percentages in parentheses indicate the amount used.
@@
Among zero-shot models, RadZero achieved the highest Dice scores on both datasets. On SIIM, it outperformed CARZero by 71% demonstrating superior segmentation capability. The smaller margin on RSNA is likely due to its coarser annotations—bounding boxes rather than pixel-level masks—which limit the advantages of RadZero’s fine-grained VL similarity maps. As in phrase grounding, the results of RadZero (224px) suggest that the superior performance is not merely driven by higher resolution.
RadZero also remained competitive against fine-tuned models. It outperformed MGCA (1%) on both datasets, demonstrating the effectiveness of zero-shot segmentation. Although it did not surpass MGCA (10%) or (100%), RadZero requires no mask labels, enabling broader generalization beyond fixed vocabularies, as further discussed in Sec. 5.3.
@@
SIIM’s detailed annotations also allow for Pix-AUC evaluation. RadZero achieved the highest score, exceeding even MGCA (10%), indicating well-calibrated VL similarity maps that distinguish positive from negative regions. In contrast, MedKLIP and CARZero, both relying on attention maps, performed worse with scores of 0.648 and 0.856, respectively. For a fair comparison, we also evaluated CARZero’s pre-softmax logits (CARZero (logits)), which improved performance but still fell short of RadZero. Notably, CARZero (logits) underperformed its own Dice score (0.081 vs. 0.100) despite threshold tuning, as expected from the inconsistent scaling of dot-product similarity. In contrast, VL-CABS directly encode pixel-level text–image similarity, allowing low values for negative samples. This contributed to its superior Pix-AUC, as further supported in Sec. 5.2.
@@
Figure 4: VL similarity maps of CXR images from CXD10, representing (a) normal, (b) fibrosis, and (c) effusion in the right lung. The value at the top-right corner represent the similarity probability $\hat{l}$ between each CXR image and the text prompt (bottom-right corner).
Figure 5: Open-vocabulary semantic segmentation results: (a), (b) for findings and (c) for anatomical regions. The CXR images and bounding box labels are from CXD10. The segmentation thresholds were set to 0.7 for (a) and (b), and 0.4 for (c).
5.2 VL similarity map analysis
@@
Figure 5 illustrates that RadZero effectively aligned visual and textual representations with VL-CABS. The outputs, VL similarity map $\hat{M}$ and probability $\hat{l}$, offer both interpretable visualizations and quantitative metrics. For the normal image Figure 5 (a), the model assigned low similarity (0.072) to the prompt “There is atelectasis,” with a dark VL similarity map, indicating weak alignment between vision tokens and the text embedding, In contrast, “There is no atelectasis” (0.921) and “The lungs are clear” (0.846) yielded bright activations across lung fields, reflecting strong alignment.
@@
Figure 5 (b) shows fibrosis, and “There is fibrosis” resulted in high similarity (0.970) with strong activations in the affected lung. Prompts indicating normality received much lower scores (0.093 and 0.014) and darker VL similarity maps, clearly distinguishing abnormal from normal descriptions.
@@
Figure 5 (c) highlights RadZero’s ability to distinguish anatomical descriptions. For right-sided pleural effusion, the model assigned high similarity (0.907) to “There is effusion,” with bright activations in the correct region. Notably, “There is right effusion” (0.946) scored even higher, indicating accurate localization, while “There is left effusion” scored much lower (0.125) and a dark VL similarity map, showing that the model correctly distinguishes between left and right lung regions.
@@
Overall, these results underscore the explainability of VL-CABS. The similarity probability is verifiable at the pixel level, enabling spatially grounded explanations. By explicitly revealing how conclusions are derived, RadZero offers enhanced interpretability in the context of disease diagnosis.
@@ 5.3 Open-vocabulary semantic segmentation
Figure 5 presents open-vocabulary semantic segmentation results for both findings and anatomical regions. Segmentation masks were generated by thresholding the VL similarity map $\hat{M}$ for each text prompt; in cases of overlapping predictions, the prompt with the highest similarity was assigned. In Figure 5 (a), RadZero successfully localized lesions based on text queries, though some segmentation masks extended beyond ground truth boxes, indicating room for improvement. Notably, certain incorrect predictions captured clinically relevant features that were not explicitly annotated: in Figure 5 (b), a chest tube was reasonably associated with “pneumothorax.” Figure 5 (c) further demonstrates RadZero’s ability to segment anatomical structures without supervision, inferring approximate spatial regions from text despite imprecise boundaries. These results highlight the potential of VL-CABS for zero-shot open-vocabulary semantic segmentation and RadZero’s capacity to align textual descriptions with medical imagery. Additional qualitative examples are provided in Appendix F.
5.4 Ablation Studies
Method Similarity Trainable layers MP Res. Classification Grounding Segmentation
OI PC PC20 CXR14 CXP CXD10 CXD10 MS-CXR SIIM
(a) dot-product Linear 224 0.839 0.824 0.853 0.805 0.896 0.792 0.472 0.784 0.078
(b) $\cos$ Linear 224 0.843 0.830 0.863 0.805 0.902 0.786 0.483 0.790 0.078
(c) $\cos$ 2 Transformer 224 0.845 0.832 0.860 0.808 0.895 0.793 0.539 0.838 0.099
(d) RadZero (224px) $\cos$ 2 Transformer 224 0.851 0.841 0.879 0.807 0.903 0.785 0.537 0.832 0.121
RadZero $\cos$ 2 Transformer 518 0.847 0.841 0.871 0.804 0.900 0.787 0.622 0.844 0.171
LiT [38] - Linear 224 0.768 0.769 0.775 0.764 0.854 0.735 - - -
dino.txt [18] - 2 Transformer 224 0.834 0.816 0.837 0.797 0.901 0.770 0.121 0.174 0.021
CARZero [19] - Transformer Dec. 224 0.827 0.815 0.877 0.795 0.889 0.770 0.437 0.743 0.072
Table 5: Ablation study of model architecture components. “MP” denotes multi-positive.
Ablation study on RadZero components.
1) Similarity function: Comparing (a) and (b) in Table 5 shows the effect of using cosine similarity instead of scaled dot-product [32] for VL alignment. Cosine similarity, which better aligns with inference-time VL similarity maps, improves classification and slightly enhances grounding. 2) Trainable parameters: (b) and (c) compare a linear layer and a two-layer Transformer, with the image encoder frozen. Transformer layers yielded consistent gains across all tasks, particularly in grounding (0.483 → 0.539) and segmentation (0.078 → 0.099). 3) Multi-positive pairs: The difference between (c) and (d) lay in the use of multi-positive contrastive pairs. (d) improved classification (e.g., PC20: 0.860 → 0.879) and segmentation (0.099 → 0.121), highlighting the advantage of richer supervision. 4) Image resolution: RadZero and (d) shared the same architecture, except that RadZero used higher resolution inputs (518 vs. 224). This change substantially improved grounding (0.537 → 0.622) and segmentation (0.121 → 0.171), showing the importance of high-resolution features for spatially localized tasks. In classification, (d) outperformed RadZero: likely due to the vision encoder (XrayDINOv2[5]) being pre-trained at 224 pixels, indicating that using an encoder pre-trained at higher resolutions may further enhance the performance.
Comparison among different VL alignment approaches.
Table 5 compares RadZero with alternative VL alignment methods, keeping all settings identical except for VL feature fusion and loss computation. LiT, which uses a [CLS] embedding for alignment, showed limited classification performance and was incapable of grounding or segmentation. dino.txt improved classification through additional Transformer layers, but its mean pooling constrained grounding and segmentation performance. CARZero introduced a cross-attention decoder, enhancing performance on those tasks. However, when compared to the RadZero ablations, (b) outperformed CARZero across most metrics, showing that VL-CABS alone is sufficiently effective.